Using the exponential dose-response model, calculate the probability of infection from drinking 2 L of water with a concentration of 3 pathogens/L from

(a) Cryptosporidium assuming a dose-response factor (r) of 0.057, and

(b) rotavirus assuming r of 0.20.
Answer #1

1 L of drinking water contains 3 pathogens, therefore 2 L water contains = 3 x 2 = 6 pathogens.

From the exponential dose-response model, the probability of infection is given by Pinf = 1- e-TD

where D= 6 pathogens

a. for cryptosporidium, r=0.057, Pinf = 1-e-0.057% Pint=1-e-0.057x6 =1-e-0.342 = 1-0.710 = 0.29

b. for rotavirus, r= 0.20,   Pint=1-e-0.20x6 =1-e-1.2 = 1-0.30 = 0.7